OpenStudy (anonymous):

solve: 5/8 x + 1/16 x = 5/16 + x the solution is x =

OpenStudy (angela210793):

is x in the denom?

OpenStudy (anonymous):

x is ater 5/8 x not on top or bottom

OpenStudy (angela210793):

tht means is on the top 5x/8-x/16=5/16+x 9x/16=5/16+x 9x/16-x=5/16 -7x/16=5/16 -7x=5 x=-5/7

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

To Lauren 123: IF your original equation reads: (5/8)x +(1/16)x= 5/16 + x ; then: Multiply both sides of your equation by 16 and you will get: 10x +1 = 5 + 16x solve this equation for x will yield -4 = 6x this will give you :-2/3 = x